# The Ultimate Configuration Management Tool? Lessons from a Mixed Methods Study of Ansible’s Challenges

### v2

@@ 4.7.1. Interviews
RQ1: We identified seven core challenge categories — Control/Data Flow and Execution, Reliability, Lack of Support, Usability, Performance and Scalability, Maintenance and Information Gaps — with Q&A emphasizing high-frequency technical issues (e.g., loops, SSH connectivity) and interviews highlighting more nuanced concerns such as stale configurations, misleading community advice, and complex setup processes.
@@ 5. RQ2: Factors Influencing Ansible Adoption
This section analyzes how interview participants first encountered Ansible and their reasons for selecting it over other IaC solutions. We go over Background and Advantages described in Section 3.3.3.
Initial Exposure to Ansible. Participants described how they discovered Ansible. The most common way was in professional settings (n = 6), where it was already an established tool. Social media platforms also played a notable role in exposure (n = 4), particularly X. As E5 recounted: “In 2016, [someone] posted a tweet about a new configuration management tool (…) I had a look.” Some participants also independently sought out Ansible (n = 3).
@@
Accessibility & Usability. A significant theme in our codebook is the role of Ansible’s accessibility and usability in driving adoption. Several participants described it as Easy to use (n = 13), Easy to onboard to new people (n = 7), and with a Low Barrier to Entry for Setup (n = 6) highlighting how they only needed to master a limited set of concepts—such as basic playbook structures and inventory configuration—before managing a wide range of environments. B1 remarked, “To get things done, I only really needed to understand a small set of concepts regarding how Ansible works (…) ”. Participants also cited Ansible’s Readable syntax (n = 6) and Good Logs (n = 1). E6 highlighted that “[we] were looking into Ansible because it has a nice YAML syntax”.
@@
Tool Qualities and Ecossystem. Beyond accessibility, Ansible was praised for its tool qualities. Participants highlighted its Efficient use of resources, with one participant citing Lightweight resource consumption and three mentioning Scalability. Efficiency was also linked to Cost-Effectiveness, with B4 saying, “Ansible is like a [inexpensive] version [of Kubernetes] for scalability”. Another advantage of Ansible is its Flexibility, since it can be a One-Tool Solution (n = 2) capable of managing heterogeneous environments (e.g., Windows, Linux, or network devices). As E3 puts it, the “One tool that does it all”. Several participants emphasized that Ansible’s Compatibility (n = 6) with various operational requirements and existing workflows set it apart from other tools, and its Agentless (n = 9) model that depends solely on SSH. Idempotence (n = 2) was also considered important for participants, ensuring the same playbook can be applied repeatedly to converge infrastructure into a consistent state. As I3 described: “Some commands are dangerous to invoke twice.” From an Ecosystem standpoint, respondents highlighted features such as Easy to extend (n = 4), Good Tooling (n = 1), and the fact that It’s implemented in Python (n = 2).
@@
Participants frequently framed their rationale in comparative terms, highlighting Ansible’s advantages over alternatives like Puppet, Chef, and Salt. A common comparative advantage cited was the elimination of agent management overhead. E5 provided an illustrative anecdote highlighting Ansible’s comparative operational model advantages: “(…) a place I was working was using Puppet, and the admin was woken up at 3 am from a page, fixed the issue, went to bed, and then Puppet restored the broken configuration, so he woke up again, fixed the Puppet, and then went to bed. Then, was woken up again because Puppet updated itself from git.”.
@@
Community. A third central theme centers on Ansible’s Community. The tool’s Popularity (n = 4) and Community (n = 4) are significant advantages. E1 noted: “Ansible’s advantages (…) are it being a well-supported IaC tool. Since there is a large community around it, it is easier to find solutions to specific problems online.” I1 summarized it as: “Ansible was just the hot topic at the time.” Participants also mentioned that Ansible’s strong market presence often led their organizations or academic institutions to mandate its use (n = 8) (“it was standard by the time I joined the team”, I3). Moreover, if it was not mandatory, it was often Recommended (n = 3) by colleagues or professors.
@@
RQ2: Ansible’s adoption is influenced by its ease of use, low barrier to entry, and agentless design. Participants highlighted how minimal setup requirements and an approachable syntax allowed them to manage diverse infrastructures quickly. They also praised Ansible’s extensibility—particularly its Python-based ecosystem. A strong community presence and widespread popularity were seen as crucial.
@@ 6. RQ3: Improving Ansible
This section describes the Recommendations described in Section 3.3.3. While participants generally acknowledged Ansible’s strengths, they also offered concrete recommendations to improve.
@@
Documentation and Learning Resources. A recurring concern expressed by participants (n = 10) was the need to improve documentation. E4 described the current official documentation as scattered, noting that it “is poor considering the other configuration tools (take a look at Puppet’s). The websites are not appealing and not properly organized. Strong concepts are scattered over different pages.”. I3 identified specific documentation challenges regarding module organization: “Documentation needs a big revamp: it’s not clear what modules are built-in or part of Ansible Galaxy.”. E8 found the documentation examples helpful but noted that “sometimes they don’t [provide] all the examples” The participants also called for better onboarding resources (n = 6) and more education (n = 9). As E3 put it “ When you educate beginners, they can identify incorrect information and avoid it.”.
@@
Debugging and Troubleshooting. Several participants highlighted limitations in Ansible’s current debugging capabilities. Better debugging tools (n = 5) and improved error messages (n = 6) were frequently mentioned. E2 stressed that “error messages when plugins fail could be improved,” as they can sometimes be “misleading”. I2 reinforced this point, saying that “more specific error messages would be very helpful. Understanding exactly where an issue presents is the most important part of the debugging process.”. Another concept frequently raised was the lack of rollback-on-failure support (n = 5).
@@
Performance and Scalability. Participant also critiqued performance and mentioned the need to improve speed (n = 7) and improve parallelization (n = 4). As E2 put “Slow runs is an eternal pain point.”. E1 advocated for integrating solutions akin to Mitogen (Mitogen Contributors, 2025), to help optimize remote execution over SSH to reduce overhead and B1 proposed a tag-based approach: “Having the ability to set group tags for tasks, and have groups of tasks be able to run in parallel would be a decent enough solution (tasks with tag A can run in parallel with tasks with tag B).”.
@@
Advanced Programming Capabilities. A group of participants called for the expansion of Ansible’s programming model. While some participants advocated for more high-level language concepts (n = 3) — including real loops, complex conditionals, and synchronization primitives — others (n = 5) stressed state management. E7, for instance, argued for “functionality in Ansible to handle state management effectively” and provided an extensive analysis of Ansible’s state management limitations, saying, “even though Ansible claims to be completely stateless, which is supposed to be its strength, I’m not so convinced. Without state management, it’s challenging to handle certain problems.”.
Participant Experience Levels. While some recommendations, such as providing more education, were popular across experience levels, other priorities varied according to participants’ familiarity with Ansible. Beginners often highlighted the need for better onboarding (n = 2) and rollback on failure (n = 3). In contrast, Experienced users (n = 4) tended to focus on improving speed and other performance enhancements, reflecting the operational demands of managing more extensive infrastructures.
@@
RQ3: Interviewees emphasized the need for better documentation, improved debugging, and enhanced error reporting. Many highlighted performance optimizations, including faster execution and parallelization, while others advocated for rollback support, stronger third-party integrations, and a more user-friendly GUI to streamline workflows.

### v3

@@ 4.1. Topic Evolution
RQ1: We identified seven core challenge categories: Control/Data Flow and Execution, Reliability, Lack of Support, Usability, Performance and Scalability, Maintenance and Information Gaps. Our Q&A analysis highlighted frequent technical issues (e.g., loops, SSH connectivity), while interviews revealed deeper concerns such as stale configurations, misleading community advice, and complex setup.
@@ 5. RQ2: Factors Influencing Ansible Adoption
This section analyzes how interview participants first encountered Ansible and the reasons they selected it over other IaC solutions.
Accessibility & Usability.
@@
A significant theme in our Advantages codebook is the role of Ansible’s accessibility in driving adoption. Several participants described it as Easy to use (n=13), Easy to onboard to new people (n=7), and with a Low Barrier to Entry for Setup (n=6) highlighting how they only needed to master a limited set of concepts—such as basic playbook structures and inventory configuration—before managing a wide range of environments. B1 remarked, “To get things done, I only really needed to understand a small set of concepts regarding how Ansible works (…)”. Participants also cited Ansible’s Readable syntax (n=6), E6 highlighted that “[we] were looking into Ansible because it has a nice YAML syntax”.
Tool Qualities & Ecosystem.
@@
Beyond accessibility, Ansible was praised for its tool qualities, with participants discussing its Scalability (n=5). Another advantage of Ansible is its Flexibility, since it can be a One-Tool Solution (n=2) capable of managing heterogeneous environments (e.g., Windows, Linux, or network devices). As E3 puts it, the “One tool that does it all”. Several participants emphasized that Ansible’s Compatibility (n=6) with various operational requirements and existing workflows set it apart from other tools, and its Agentless (n=9) model that depends solely on SSH. Idempotence (n=2) was also considered important for participants, ensuring the same playbook can be applied repeatedly to converge infrastructure into a consistent state. As I3 described: “Some commands are dangerous to invoke twice.” From an Ecosystem standpoint, respondents highlighted features such as Easy to extend (n=4), and the fact that It’s implemented in Python (n=2).
@@
Participants frequently framed their rationale in comparative terms, highlighting Ansible’s advantages over alternatives like Puppet, Chef, and Salt. A common comparative advantage cited was the elimination of agent management overhead.
E5 provided an illustrative anecdote highlighting Ansible’s comparative operational model advantages: “(…) a place I was working was using Puppet, and the admin was woken up at 3 am from a page, fixed the issue, went to bed, and then Puppet restored the broken configuration, so he woke up again, fixed the Puppet, and then went to bed. Then, was woken up again because Puppet updated itself from git.”
Community.
@@
A third central theme centers on Ansible’s Community. The tool’s Popularity (n=4) and Community (n=4) are significant advantages. E1 noted: “Ansible’s advantages (…) are it being a well-supported IaC tool. Since there is a large community around it, it is easier to find solutions to specific problems online.” I1 summarized it as: “Ansible was just the hot topic at the time.” Participants also mentioned that Ansible’s strong market presence often led their organizations to mandate its use (n=8) (“it was standard by the time I joined the team”, I3). Moreover, if it was not mandatory, it was often Recommended (n=3) by colleagues or professors.
@@
RQ2: Ansible’s adoption is influenced by its ease of use, low barrier to entry, and agentless design. Participants highlighted that minimal setup requirements and an approachable syntax enabled them to quickly manage diverse infrastructures. They also praised Ansible’s extensibility, strong community presence, and popularity.
@@ 6. RQ3: Improving Ansible
This section describes the Recommendations described in Section 3.3. While participants generally acknowledged Ansible’s strengths, they also offered concrete recommendations to improve.
Documentation and Learning Resources.
@@
A recurring concern expressed by participants (n=10) was the need to improve documentation. E4 described the current official documentation as scattered, noting that it “is poor considering the other configuration tools (take a look at Puppet’s). The websites are not appealing and not properly organized”. Strong concepts are scattered over different pages.” I3 identified specific documentation challenges regarding module organization: “Documentation needs a big revamp: it’s not clear what modules are built-in or part of Ansible Galaxy.” E8 found the documentation examples helpful but noted that “sometimes they don’t [provide] all the examples” The participants also called for better onboarding resources (n=6) and more education (n=9). As E3 put it “ When you educate beginners, they can identify incorrect information and avoid it.”
@@
Debugging and Troubleshooting. Several participants highlighted limitations in Ansible’s current debugging capabilities. Better debugging tools (n=5) and improved error messages (n=6) were frequently mentioned. E2 stressed that “error messages when plugins fail could be improved,” as they can sometimes be “misleading”. I2 reinforced this point, saying that “more specific error messages would be very helpful. Understanding exactly where an issue presents is the most important part of the debugging process”. Another concept frequently raised was the lack of rollback-on-failure support (n=5).
Performance and Scalability.
@@
Participants also critiqued performance and mentioned the need to improve speed (n=7) and improve parallelization (n=4). As E2 put “Slow runs is an eternal pain point.” E1 advocated for integrating solutions akin to Mitogen (Mitogen Contributors, 2025), to help optimize remote execution over SSH to reduce overhead and B1 proposed a tag-based approach: “Having the ability to set group tags for tasks, and have groups of tasks be able to run in parallel would be a decent enough solution (tasks with tag A can run in parallel with tasks with tag B).”
Advanced Programming Capabilities.
@@
A group of participants called for the expansion of Ansible’s programming model. While some participants advocated for more high-level language concepts (n=3) (e.g., real loops, complex conditionals, and synchronization primitives), others (n=5) stressed state management. E7, for instance, argued for “functionality in Ansible to handle state management effectively” and provided an extensive analysis of Ansible’s state management limitations, saying, “even though Ansible claims to be completely stateless, which is supposed to be its strength, I’m not so convinced. Without state management, it’s challenging to handle certain problems.”
@@
RQ3: Interviewees emphasized the need for better documentation, improved debugging, and enhanced error reporting. Many highlighted performance optimizations, including faster execution and parallelization, while others advocated for rollback support and stronger third-party integrations.